We aligned the PSL green optics so that the PSL green beam and Y arm green beam interfere. 2 beams are now hitting the Y arm beat PD. The DC level from the beat PD is about 13 mV.
We didn't try to see the beat signal for today, because the temperature of the doubling crystal seemed funny. We need to look into it tommorow.
Currently, the temperature control is enabled and the set point is 36.9 deg C, but the temperature is stuck at 33.0 deg C.